Find the confidence interval specified.
The mean score,x , on an aptitude test for a random sample of 5 students was 73. Assuming that σ=15 construct a 95.44% confidence interval for the mean score, μ of all students taking the test.

Answer by stanbon(75887)

Find the confidence interval specified.
The mean score,x , on an aptitude test for a random sample of 5 students was 73. Assuming that σ=15 construct a 95.44% confidence interval for the mean score, μ of all students taking the test.
-------
(1-0.9544)/2 = 0.0228
=====
-invNorm(0.0228) = 1.9991
--------------
x-bar = 73
ME = z*s/sqrt(n) = 1.9991*(15/sqrt(73) = 3.51
-----
95.44% CI:: 73-3.51 < u < 73+3.51
---------------
Cheers,
Stan H.
--------------